# LEO-VL: Towards 3D Vision-Language Generalists via Data Scaling with Efficient Representation

### v1

@@ Abstract
Developing 3D-VL generalists capable of understanding 3D scenes and following natural language instructions to perform a wide range of tasks has been a longstanding goal in the 3D-VL community. Despite recent progress, 3D-VL models still lag behind their 2D counterparts in capability and robustness, falling short of the generalist standard. A key obstacle to developing 3D-VL generalists lies in data scalability, hindered by the lack of an efficient scene representation. We propose LEO-VL, a 3D-VL model built upon condensed feature grid (CFG), an efficient scene representation that bridges 2D perception and 3D spatial structure while significantly reducing token overhead. This efficiency unlocks large-scale training towards 3D-VL generalist, for which we curate over 700k high-quality 3D-VL data spanning four domains of real-world indoor scenes and five tasks such as captioning and dialogue. LEO-VL achieves state-of-the-art performance on a variety of 3D QA benchmarks, including SQA3D, MSQA, and Beacon3D. Ablation studies confirm the efficiency of our representation, the importance of task and scene diversity, and the validity of our data curation principle. Furthermore, we introduce SceneDPO, a novel post-training objective that enhances the robustness of 3D-VL models. We hope our findings contribute to the advancement of scalable and robust 3D-VL generalists.
![[Uncaptioned image]](x1.png)
@@ 1 Introduction
With the crucial role of 3D scene understanding in human intelligence, building generalist models that can understand 3D environments and communicate through natural language [34, 97, 78, 55] has become a central goal of the 3D vision-language (3D-VL) community. Recent progress has advanced capabilities across 3D-VL tasks such as grounding [117, 118, 39], captioning [53, 93], reasoning [32, 46], and dialogue [34, 95]. Nonetheless, current 3D-VL models fall significantly short of the generalist capabilities demonstrated by their 2D counterparts [60, 82, 88, 13, 42]. We posit that a key obstacle to developing 3D-VL generalists lies in data scalability, hindered by the absence of an efficient scene representation. Therefore, establishing an efficient and scalable scene representation is a critical first step towards building 3D-VL generalists.
@@
Existing scene representations in 3D-VL models can be categorized into two approaches. The first adopts 3D modality inputs such as point clouds, but requires complex pre-processing pipelines such as 3D reconstruction and instance segmentation [117, 39, 11]. And they confront the inherent challenge of direct 3D perception, which is exacerbated by the scarcity of 3D-VL data. The other approach leverages 2D vision-language (2D-VL) models for perception [115, 111, 71], fueled by the advancement of 2D perception. However, this approach suffers from poor data scalability, as representing a 3D scene with multi-view 2D features incurs significant computational overhead. Additionally, both approaches typically use rudimentary 3D position embeddings for 3D spatial modeling, which fall short in capturing 3D spatial structures effectively.
We propose LEO-VL, a 3D-VL model built on condensed feature grid (CFG), an efficient scene representation that bridges 2D perception and 3D spatial structure, significantly improving efficiency while preserving the strength of 2D perception. Specifically, we employ a pretrained 2D visual encoder to extract image features, which are then back-projected into 3D space for voxelization. We utilize rotary position embedding (RoPE) [79] to encode the height of voxels and condense voxel features vertically to form a CFG. For horizontal spatial modeling, we integrate 2D Fourier features [80] into the CFG. In addition to efficiency, our CFG simplifies 3D spatial modeling by disentangling vertical and horizontal dimensions. The scene tokens of CFG are further connected to a large language model (LLM) for auto-regressive language modeling.
For the training of LEO-VL, we incorporate five 3D-VL tasks, including captioning, question answering (QA), planning, and dialogue, spanning four domains of real-world indoor scenes: ScanNet [19], 3RScan [86], MultiScan [58], and ARKitScenes [5]. Within each domain, we curate existing datasets and newly generated data while prioritizing data quality over scale. This could avoid model performance degradation from naive scaling with low-quality sources. With our curation principle, we collect a balanced and diverse corpus of over 700k high-quality 3D-VL data samples for the instruction tuning of LEO-VL.
We evaluate LEO-VL on a variety of 3D QA benchmarks, demonstrating state-of-the-art performance with significantly higher efficiency than prior methods. Ablation studies exhibit advantages of the CFG in efficiency and 3D spatial modeling. Ablations on various training data configurations reveal the benefits of tasks such as captioning and dialogue, and the importance of scene diversity for strong cross-domain performance. We further show that scaling with simplistic QA data leads to performance degradation, confirming our data curation principle, while our curated data yields consistent scaling effects. Additionally, we introduce SceneDPO, a novel post-training objective that incorporates contrasts across answers and scenes, and demonstrate its effectiveness in enhancing model robustness compared to conventional supervised finetuning (SFT).
@@
We propose LEO-VL, a 3D-VL generalist built upon condensed feature grid (CFG), an efficient scene representation that enables strong capability in 3D scene perception with significantly improved efficiency and simplified 3D spatial modeling.
@@
We construct a comprehensive 3D-VL instruction-tuning dataset spanning four scene domains and five tasks. Our dataset comprises over 700k high-quality samples, striking a balance between scale and quality, and serving as a reliable foundation for the training of 3D-VL generalists.
@@
We demonstrate LEO-VL’s state-of-the-art performance across 3D QA benchmarks such as SQA3D. Ablation studies validate the efficiency of our representation, the benefits of task and domain diversity, and confirm our data curation principle. We also showcase the superiority of our post-training objective in enhancing model robustness compared to conventional SFT.
@@ 2 Related Work
Early work in 3D-VL understanding develops models that process 3D representations such as point clouds [67, 69, 65, 45, 91, 68, 108, 37, 100] and voxels [27, 59, 74, 81, 28, 18, 15, 76]. With the maturation of pretraining and instruction tuning techniques, 3D-VL models have demonstrated significant performance improvements [109, 36, 1, 10, 36, 49, 94, 116, 117, 114, 90] and the capability of generalist models [29, 93, 34, 95, 11, 70, 118, 26, 107, 16, 32, 21]. Nonetheless, 3D perception remains challenging due to the data scarcity and inherent complexity of 3D representation space. In contrast, an alternative line of work leverages 2D perception models to handle 3D-VL tasks [64, 77, 23, 38, 31, 24, 57, 52], a strategy made increasingly effective by advances in 2D vision-language models [48, 88, 106, 115, 111, 71]. However, adapting 2D VLMs for 3D scene understanding poses two challenges: the cost of representing 3D scenes and the absence of effective 3D spatial modeling. While recent efforts [113, 99, 83, 33] attempt to address these issues, they still fall short of achieving both high efficiency and strong capability. In this work, we propose a framework that balances efficiency and effectiveness, showing the potential of foundational 3D-VL generalist models.
@@
3D-VL datasets and benchmarks are typically built upon 3D scene assets [19, 98, 86, 58, 5, 7, 73, 8, 41, 112, 25, 20]. Existing 3D-VL datasets primarily focus on object grounding [9, 2, 105] and QA [3, 56] tasks, but most of them are centered around ScanNet [19]. As the 3D-VL community advances, recent efforts have focused on aggregating diverse scene domains to enable large-scale pretraining [117, 118, 39, 89] and generalist instruction tuning [34]. This further drives the development of cross-domain datasets [96, 54, 46, 78, 102] and benchmarks [46, 35, 97]. Following this trend, our simple yet efficient framework is the first to scale 3D-VL generalist learning to the SceneVerse level [39], covering four scene domains [19, 86, 58, 5] and five 3D-VL tasks. We further demonstrate its effectiveness through comprehensive evaluations across these heterogeneous 3D scenes.
Vision-Language Learning Objective.
Prior studies [22, 35, 63, 104] indicate that SFT, as a common learning objective for 3D-VL models, may undergo weak robustness given the scarce data and overfitting risk in 3D-VL learning. Existing efforts to remedy the robustness mainly involve data augmentation [96, 40, 110], with the learning objective underexplored. Motivated by the success of reinforcement learning from human feedback (RLHF) [61, 75, 72, 101, 62, 50, 17] for post-training, prior 2D-VL models [43, 66, 87, 92] employ direct preference optimization (DPO) objective [72] to address the robustness issue. We bring these insights into 3D-VL learning and introduce an additional objective with scene contrast to post-training for more robust 3D-VL models.
@@ 3 Method
In this section, we outline our approach to 3D-VL generalists. In Section 3.1, we introduce our model, LEO-VL, and particularly its efficient representation of 3D scenes, which largely reduces the token overhead while preserving strong perception capability. In Section 3.2, we introduce the recipe of our comprehensive instruction-tuning datasets for training of 3D-VL generalists. In Section 3.3, we further introduce a novel post-training objective to enhance the robustness of 3D-VL models.
@@ 3.1 Model
As illustrated in Fig. 2, LEO-VL extracts image features from RGBD frames and transforms the features into an efficient scene representation, termed the condensed feature grid (CFG). Next, an LLM performs auto-regressive language modeling over the joint sequence of CFG tokens and instruction tokens. In contrast to prior 3D-VL models that rely on object-centric 3D representation [34, 118, 32, 46], LEO-VL adopts 2D perception considering the limitations of the object-centric approach, such as the dependency of instance segmentation and loss of scene-level information.
@@
We employ a 2D visual encoder to extract image features from multi-view RGB inputs. Given a set of RGB images , the encoder produces corresponding feature maps , where , , and denote the resolution and feature dimension. For each view, the corresponding depth maps are used to back-project 2D pixels into 3D space. Let and denote the intrinsic and extrinsic camera matrices, respectively. A pixel at coordinate with depth is back-projected to a 3D point through the following homogeneous transformation:
We compute 3D coordinates for all pixels and interpolate them to align with the feature map resolution . These 3D points are then associated with corresponding 2D features.
@@
First, we voxelize the 3D points and average pool the point features within each voxel, discarding empty voxels. For referring tasks such as object captioning, a learnable anchor embedding is injected into voxels that fall within the referred region. Despite voxelization, the representation remains costly and still limits scalability. To further improve efficiency, we convert voxels into the CFG by pooling voxel features along the height axis. Before pooling, we encode the height of each voxel with RoPE [79]. To encode horizontal positions, we integrate 2D Fourier features [80] in the CFG.
Position Embeddings.
We adopt RoPE for vertical spatial modeling and Fourier features for horizontal spatial modeling. RoPE encodes position information by rotating the feature vector by a series of angles for different feature channels. The angles are derived from linear combinations of the position value and channel-specific frequencies. The rotary nature enables RoPE to better capture vertical spatial relations, such as distinguishing objects at different heights, compared to additive position embeddings [85, 80]. In contrast, the Fourier feature works by first applying a linear transformation to the position value and then computing sinusoidal components, which are passed through an MLP to produce embeddings. Let denote the rotation matrix for RoPE and the projection matrix for the Fourier feature. The formulations of position-encoded features are as follows:
@@ 3.2 Data
The construction of LEO-VL training data is guided by three principles: (1) the diversity in scene domains, which demonstrates crucial for 3D-VL learning [39]; (2) the diversity of tasks, which we think necessary given the benefits of long-response tasks [34, 54]; and (3) the balance between quality and scale, considering the potential risks in hindering model performance from low-quality data. We detail the composition of LEO-VL training data in Table 1.
@@
We include four domains of real-world indoor scenes: ScanNet [19], 3RScan [86], MultiScan [58], and ARKitScenes [5], following prior practices in scaling 3D scenes for 3D-VL learning [39, 46]. We exclude scene domains that lack attribute-rich scene graphs (e.g., HM3D [73]), which are essential for LLM-assisted data generation. We follow the official splits for ScanNet and 3RScan, SceneVerse splits [39] for MultiScan, and MSQA splits [46] for ARKitScenes.
@@
We incorporate five prevalent 3D-VL tasks with natural language outputs, all compatible with a unified instruction tuning framework. We exclude the 3D object grounding task due to its different formulation. We think the absence of the grounding task is not detrimental given the success of 2D-VL generalists without explicit grounding supervision [47, 60, 88, 13]. In addition to existing datasets, we also generate new data by prompting LLMs with scene graphs [34, 46]. Our data covers five categories of text-output tasks:
Object Captioning. This task requires the model to describe a specific object in natural language. We adopt object captions from SceneVerse [39] and MMScan [54]. We exclude datasets that repurpose object grounding texts as captions [14, 2], as such texts lack diverse descriptions regarding object attributes.
Planning. Given a high-level goal (e.g., “organize the room”), the model is required to generate a step-by-step grounded plan. We use the planning data from LEO [34] for 3RScan, and generate new data for other scene domains.
Dialogue. This task requires the model to generate context-aware responses conditioned on both the 3D scene and dialogue history. We use the dialogue data from LEO [34] for 3RScan, and generate new data for other scene domains.

### v2

# LEO-VL: Efficient Scene Representation for Scalable 3D Vision-Language Learning

@@ Abstract
Developing vision-language models capable of understanding 3D scenes has been a longstanding goal in the 3D-VL community. Despite recent progress, 3D VLMs still fall short of their 2D counterparts in capability and robustness. A key bottleneck is that current scene representations struggle to balance performance and efficiency: competitive performance comes at the cost of heavy token overhead, which in turn hampers the scalability of 3D-VL learning. To address this, we propose the condensed feature grid (CFG), an efficient scene representation featuring significantly reduced token overhead and strong perception capability. Building on CFG, we introduce LEO-VL, a 3D VLM trained on 700k 3D-VL data spanning four real-world indoor domains and five tasks such as captioning and dialogue. To enhance the robustness of 3D VLM, we further propose SceneDPO for post-training, which involves contrasts across answers and scenes. LEO-VL achieves state-of-the-art performance on various 3D QA benchmarks, including SQA3D, MSQA, and Beacon3D. Our extensive experiments highlight the efficiency of our representation, the benefit of task and scene diversity, consistent scaling effects, and the advantages of SceneDPO compared to SFT and GRPO. We hope our findings advance the efficiency, scalability, and robustness of future 3D VLMs. ††Corresponding author.
@@ 1 Introduction
With the crucial role of 3D scene understanding in human intelligence, building 3D VLMs that can understand 3D environments and communicate with humans through natural language has become a central goal of the 3D-VL community (Huang et al., 2024b; Yang et al., 2025b; Song et al., 2025; Ma et al., 2024). Recent progress has advanced capabilities across 3D-VL tasks such as grounding (Zhu et al., 2023; 2024b; Jia et al., 2024), captioning (Luo et al., 2023; Xu et al., 2024), reasoning (Huang et al., 2024a; Linghu et al., 2024), and dialogue (Huang et al., 2024b; Yang et al., 2024). Nonetheless, current 3D VLMs fall significantly short of the capabilities exhibited by their 2D counterparts (OpenAI, 2023; Team et al., 2023; Wang et al., 2024b; Chen et al., 2024c; Li et al., 2024a). We argue that a key obstacle to advancing 3D VLMs lies in scene representation, which struggles in the performance-efficiency trade-off and could hinder the scalability of 3D-VL learning.
@@
Existing scene representations for 3D VLMs can be categorized into two approaches. The first adopts 3D modality input such as point cloud, requiring complex pre-processing pipelines such as 3D reconstruction and instance segmentation (Zhu et al., 2023; Jia et al., 2024). This approach also confronts the inherent difficulty of 3D perception, which is challenging given the scarcity of 3D-VL data. The second approach uses 2D images or videos as input, leveraging the strength of 2D perception (Zhu et al., 2024a; Zheng et al., 2025). Despite stronger performance, it incurs substantial token overhead and severely limits data scalability. Additionally, both approaches typically adopt rudimentary 3D position embeddings, which may fall short in modeling complex spatial structures.
To overcome these limitations, we propose the condensed feature grid (CFG), a novel scene representation that significantly improves efficiency while preserving strong perception capability. Given multi-view RGB-D inputs, we back-project 2D features to 3D voxels, encode their height features using rotary position embedding (RoPE) (Su et al., 2024), and vertically condense voxels within each pillar into grid tokens. We further inject horizontal position embeddings with 2D Fourier features (Tancik et al., 2020). This design reduces scene token overhead to only , while maintaining accurate 3D structures. Based on this representation, we build our model LEO-VL by integrating the CFG tokens with a large language model (LLM) for auto-regressive language modeling.
As the representation efficiency unlocks scalability of 3D-VL learning, we construct a comprehensive 3D-VL dataset spanning four real-world indoor domains (Dai et al., 2017; Wald et al., 2019; Mao et al., 2022; Baruch et al., 2021) and five 3D-VL tasks. We curate existing data and generate new samples where necessary. We emphasize data quality over scale to avoid model performance degradation, and ultimately collect over 700k diverse and high-quality data samples for 3D-VL instruction tuning.
Notably, prior studies suggest that supervised finetuning (SFT) alone can be insufficient to build robust 3D VLMs given the prevalent overfitting risks (Deng et al., 2024; Huang et al., 2025b). To enhance the robustness of 3D VLMs, we propose SceneDPO, a post-training objective that involves contrasts across answers and scenes. We further demonstrate that SceneDPO yields better in-domain and out-of-domain performances compared to SFT and group relative policy optimization (GRPO).
We show that LEO-VL achieves state-of-the-art performance with significantly higher efficiency on various 3D QA benchmarks, including ScanQA, SQA3D, MSQA, and Beacon3D. Ablation studies highlight the efficiency and spatial modeling of CFG, the impact of task diversity such as captioning and dialogue, and the importance of scene diversity for strong cross-domain performance. Moreover, we observe that naive scaling with low-quality QA data degrades performance, whereas our data curation principle yields consistent scaling effects.
@@
We propose LEO-VL, a 3D VLM built on condensed feature grid (CFG), an efficient scene representation featuring strong perception capability and simplified 3D spatial modeling.
@@
We construct a comprehensive 3D-VL instruction-tuning dataset of 700k samples across four domains and five tasks, prioritizing data quality to ensure reliable scaling for 3D-VL learning.
@@
We demonstrate LEO-VL’s state-of-the-art performance across various 3D QA benchmarks, supported by extensive ablations on representation efficiency, task and domain diversity, and data quality. We further introduce SceneDPO as an effective post-training objective for 3D VLM.
@@ 2 Related Work
Most early works in 3D-VL understanding develop models based on 3D representations such as point clouds (Qi et al., 2017a; b; Phan et al., 2018; Li et al., 2018; Wu et al., 2019; Qi et al., 2019; Zhao et al., 2021a; Huang et al., 2021; Yu et al., 2022) and voxels (Graham, 2015; Maturana & Scherer, 2015; Riegler et al., 2017; Tatarchenko et al., 2017; Graham et al., 2018; Dai & Nießner, 2018; Choy et al., 2019; Schult et al., 2023). With the maturation of 3D-VL pretraining and instruction tuning techniques, 3D VLMs have demonstrated significant improvements in capabilities (Zhao et al., 2021b; Huang et al., 2022; Abdelreheem et al., 2022; Chen et al., 2022; Huang et al., 2022; Liu et al., 2023b; Xue et al., 2024; Zhu et al., 2025; 2023; Zhou et al., 2024; Wang et al., 2025; Hong et al., 2023; Xu et al., 2024; Huang et al., 2024b; Yang et al., 2024; Chen et al., 2024a; Qi et al., 2024; Zhu et al., 2024b; Fu et al., 2025; Zhang et al., 2024b; Chu et al., 2024; Huang et al., 2024a; Deng et al., 2025). In contrast, an alternative line of work leverages 2D perception models to handle 3D-VL tasks and exhibits strong performances (Peng et al., 2023; Shafiullah et al., 2023; Ding et al., 2023; Jatavallabhula et al., 2023; Huang et al., 2023; El Banani et al., 2024; Man et al., 2024; Luo et al., 2025; Zhu et al., 2024a; Zheng et al., 2025; Qi et al., 2025). However, adapting 2D VLMs to 3D scene understanding poses two challenges: the cost of representing 3D scenes and the absence of effective 3D spatial modeling. Despite recent efforts (Zhi et al., 2025; Yu et al., 2025; Thai et al., 2025; Huang et al., 2025a), the scene representation still struggles at the performance-efficiency trade-off. In this paper, we propose a novel scene representation that achieves both strong performance and high efficiency, and introduce a new perspective in 3D spatial modeling.
@@
Based on indoor 3D scene assets (Dai et al., 2017; Yeshwanth et al., 2023; Wald et al., 2019; Mao et al., 2022; Baruch et al., 2021; Chang et al., 2017; Ramakrishnan et al., 2021; Chang et al., 2015; Khanna et al., 2024; Zheng et al., 2020; Fu et al., 2021; Deitke et al., 2022), existing 3D-VL datasets primarily focus on object grounding (Chen et al., 2020; Achlioptas et al., 2020; Zhang et al., 2023) and question answering (QA) (Azuma et al., 2022; Ma et al., 2023) tasks. As the 3D-VL community advances, recent efforts have focused on aggregating diverse scene domains to enable large-scale pretraining (Zhu et al., 2023; 2024b; Jia et al., 2024; Wang et al., 2024c) and instruction tuning (Huang et al., 2024b), followed by more cross-domain datasets (Yang et al., 2025a; Lyu et al., 2024; Linghu et al., 2024; Song et al., 2025; Zhang et al., 2025a) and benchmarks (Linghu et al., 2024; Huang et al., 2025b; Yang et al., 2025b). Given the potential of cross-domain scaling for 3D-VL learning, we compile a comprehensive 3D-VL data scheme covering four real-world scene domains and five instruction-tuning tasks. We further demonstrate the benefits of domain and task diversity through comprehensive evaluations.
Post-training for Vision-Language Models.
Prior studies (Deng et al., 2024; Huang et al., 2025b; Peng et al., 2025; Zhang et al., 2025b) suggest that SFT, as a common strategy for 3D VLMs, may undergo weak robustness given the scarce data and overfitting risk in 3D-VL learning. To address this, existing efforts mainly involve data augmentation (Yang et al., 2025a; Kang et al., 2024; Zhao et al., 2024), with the learning objective underexplored. Motivated by the success of reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022; Schulman et al., 2017; Rafailov et al., 2023; Yuan et al., 2024; Pang et al., 2024; Liu et al., 2024; Chu et al., 2025), recent works (Li et al., 2024b; Pi et al., 2024; Wang et al., 2024a; Xie et al., 2024) demonstrate the efficacy of direct preference optimization (DPO) (Rafailov et al., 2023) on 2D VLMs. In contrast, we introduce SceneDPO as a novel and effective post-training objective tailored for 3D VLM.
@@ 3 Method
In Section˜3.1, we introduce the model design of LEO-VL, which features an efficient representation of 3D scenes. In Section˜3.2, we introduce the recipe of our comprehensive 3D-VL instruction-tuning datasets, which spans four real-world indoor domains and five tasks. In Section˜3.3, we further introduce a novel post-training objective to enhance the robustness of 3D VLMs.
@@ 3.1 Model
Overview.
As illustrated in Fig.˜2, given multi-view RGB-D images, LEO-VL extracts 2D visual features and transforms them into the condensed feature grid (CFG). Next, an LLM performs auto-regressive language modeling over the joint sequence of CFG tokens and language tokens. In contrast to prior object-centric 3D VLMs (Huang et al., 2024b; Zhu et al., 2024b; Huang et al., 2024a; Linghu et al., 2024), LEO-VL employs 2D perception to address limitations such as the dependency of object masks, scene-level information loss, and poor perception capability. In contrast to recent video-based VLMs (Zhu et al., 2024a; Zheng et al., 2025; Qi et al., 2025), LEO-VL substantially reduces the representation overhead and simplifies 3D spatial modeling.
@@
We employ a 2D visual encoder to extract image features from multi-view RGB inputs. Given a set of RGB images , the encoder produces feature maps , where , , and denote the resolution and feature dimension. The corresponding depth maps are interpolated to patch resolution and then back-projected into 3D space. These 3D points will be assigned 2D features . Let and denote the intrinsic and extrinsic camera matrices, respectively. For example, a pixel at coordinate with depth is back-projected to a 3D point as follows:
Voxelization.
We first voxelize the 3D points by averaging point features within each voxel, with empty voxels discarded. For referring tasks such as object captioning, a learnable anchor embedding is added to voxels that fall within the referred region. This yields an initial voxel-based representation , which remains computationally expensive and poses challenges for 3D spatial modeling.
@@
To further improve efficiency, our core is to condense voxels into a 2D planar grid by pooling voxel features vertically within each pillar. To address the challenge of 3D spatial modeling, we disentangle the position encoding along vertical and horizontal directions, respectively. The process of converting voxels into CFG consists of three stages.
Vertical Position Encoding. We encode the vertical position (i.e., height) of each voxel using RoPE (Su et al., 2024), which applies a rotation-like transformation to the feature. The rotation angles relate to a pre-defined channel-wise frequencies and position value (i.e., height). Compared to additive position embeddings (Vaswani et al., 2017; Tancik et al., 2020), the rotary nature enables RoPE to better capture vertical spatial relations, such as distinguishing objects at different heights (see example in Appendix). The height-encoded voxel feature is formulated as:
Vertical Condensation. Let denote the count of voxels located in the pillar at . For each horizontal position where , the conversion from voxels to CFG token is formulated as:
Horizontal Position Encoding. We encode the horizontal position with 2D Fourier features (Tancik et al., 2020), which applies a linear transformation (denoted by ) to the position value and then computes sinusoidal components, followed by an MLP to enhance expressiveness. Let denote the horizontal position, the final CFG token is formulated as:
@@ 3.2 Data
The construction of LEO-VL training data is guided by three principles: (1) the diversity of scene domains, which demonstrates crucial for 3D-VL learning (Jia et al., 2024); (2) the diversity of tasks, which we think necessary given the benefits of long-response tasks (Huang et al., 2024b; Lyu et al., 2024); and (3) the balance between quality and scale, considering the potential risks in hindering model performance from low-quality data. We detail the composition of our data in Table˜1.
@@
We include four real-world indoor domains: ScanNet (Dai et al., 2017), 3RScan (Wald et al., 2019), MultiScan (Mao et al., 2022), and ARKitScenes (Baruch et al., 2021), following prior practices in scaling 3D scenes (Jia et al., 2024; Linghu et al., 2024). We exclude scene domains that lack attribute-rich scene graphs (e.g., HM3D (Ramakrishnan et al., 2021)), which are essential for LLM-assisted data generation.
@@
We incorporate five prevalent 3D-VL tasks with natural language outputs, which are compatible with a unified instruction tuning framework. We exclude the 3D object grounding task due to its different formulation. We think the absence of the grounding task is not detrimental, given the success of 2D VLMs without explicit grounding supervision (Liu et al., 2023a; OpenAI, 2023; Wang et al., 2024b; Chen et al., 2024c). In addition to existing datasets, we generate new data by prompting LLMs with scene graphs (Huang et al., 2024b; Linghu et al., 2024) to fill the blank (“✓” entries) in Table˜1. Our data covers five categories of text-output tasks:
Object Captioning. This task is to describe a specific object in natural language. We adopt object captions from SceneVerse (Jia et al., 2024) and MMScan (Lyu et al., 2024). We exclude datasets that re-purpose object grounding texts as captions (Chen et al., 2021; Achlioptas et al., 2020), as they lack diverse descriptions regarding object attributes.
Planning. This task is to generate a step-by-step grounded plan for a high-level goal (e.g., “organize the room”). We use the planning data from LEO (Huang et al., 2024b) for 3RScan, and generate new data for other scene domains.
Dialogue. This task concerns generating responses conditioned on both the 3D scene and dialogue context. We use the dialogue data from LEO (Huang et al., 2024b) for 3RScan, and generate new data for other scene domains.